LBTO Technical (D if time/weather permitting)

OSA: G. Bechetti

ISp: E. Solheid

ISA: D. Thompson, B. Rothberg (Tucson)

Manager: J. Urban

=============================

UT 02:02 - Open and attempting MOS acquisition test - having issues with MOS and Mask in FPU

UT 02:05 - MOS seems confused, had to manually put a mask in FPU, but when re-ran script MOS put the mask back into storage, complaining then no mask in FPU

UT 02:12 - XML file only lists 900946 not 900946.NGC7078C - could be an xml error

UT 02:13 - running modified XML file with correct iD "900946.NGC7078C"

UT 02:20 - missed an instance in the XML file - LUCI2 software put mask back into storage

UT 02:22 - Resending modified (v2) XML

UT 02:34 - Stopped at MOS acquisition - unable to find PIXSCALE in header

UT 02:44 - Error is due to typo in MASK list, a '.' was left out between '91' and 'NGC'

UT 02:53 - error "FAILED TO FIND/GET FITS-HEADER KEYWORD: PIXSCALE EXCEPTION: NULL" - could be that mask list didn't update and save

UT 02:55 - Trying MOS acquisition again

UT 03:02 - Still getting same error - restarting MOS unit, then will reinitialize and try again

UT 03:04 - Initializing MOS unit

UT 03:04 - re-executing script

UT 03:14 - still an issue with the LMS name

UT 03:15 - Shutting down and restarting LUCI2

UT 03:27 - Astrometry shows up in Aladdin, new error: "Task: failed to draw LMS reference targets. Message: No matching LMS- targets coordinates found. Exception: Null" - no longer got "PIXSCALE" error and no more "GEIRS error"

UT 03:35 - PIXSCALE is now back in the building.

UT 03:40 - Turned over scope to J.Hill and D.Miller

UT 06:10 - Taking data on transforms - using current transform in place - taking new data with BS9107

UT 07:08 - Running idl routine transform_collect - Stone_A3.list

UT 07:11 - First image luci2.20151027.0021.fits (sky)

-->#20,23 failed

UT 07:30 - restated on start=7 or GS=11

-->#26 failed

UT 07:35 - restarted on start=9 or GS=22

-->#22 failed

UT 08:23 - Redid list now using StoneD (Stone_D_new_Oct27.list)

UT 08:27 - first image luci2.20151027.0028.fits (on axis star) - second image also blank

UT 10:23 - GS#30 failed, restarted to do last one GS#20

UT 10:30 - Completed transform

===================================================

UT 10:31 - Moving to V380 Orionis - Airmas 1.293

UT 10:40 - luci2.20151027.0065 - 67 - acquisition for V380 Orionis

UT 10:52 - seems to have inserted two extra thru-slit images - 72,73

UT 11:01 - Taking data, airmass 1.303 seeing 0.8"

UT 11:04 - Aborting the script - something isn't right with the dithering - #076 was readout

UT 11:07 - Abort didn't work (reloaded and paramters were still set with grating in place)

UT 11:09 - Reloaded again, this time mirror moved into place but slit is still in FPU

UT 11:20 - Looks like 4 extra thru-slits inserted via Scriptor - will have to skip to grating set up

UT 11:24 - #0080 is start of new set

UT 11:33 - #0084 first spectroscopic exposure seeing 0.6" airmass 1.340

UT 11:53 - #0093 last V380 Orionis

UT 11:55 - Moving to Telluric HD 34481

UT 12:05 - #0098 first spectroscopic exposure for HD 34481 Telluric Star - last #101

UT 12:13 - Moving to HD 74721 - Telluric for 1.5"slit HK

UT 12:15 - Scriptor issue - the PA was 307 but apparently the guidestar selected was for PA=0 - guidestar out of patrol field error - attempted another target and scheduler vanished from GUI!

UT 12:21 - LUCI-2 responsive to filter and mask moves but scheduler is empty

UT 12:24 - resent a preset -but an active preset myight be there

UT 12:27 - LUCI-2 seems confused again - mask is in FPU when should be in turnout

UT 12:34 - reloaded script, collimation completed, nothing is happening after this step

UT 12:53 - Stopped before taking spectra for telluric - closing dome

UT 13:04 - LUCI-2 cal unit IN

UT 13:05 - zJ 200 0.5" Halo 2 flats

UT 13:19 - network is slowing down LUCI-2 GUIs, but counts seem about right for zJ flats

UT 13:45 - everytime a lamp is turned on/off the whole GUI freezes

UT 13:58 - finished cals - skipped zJ Xenon. blinds in, cal lamp out, mirror in, 0.5" in FPU

-- %USERSIG{BarryRothberg - 2015-10-27}%

Comments


 
Topic revision: r4 - 27 Oct 2015, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback